Write a Python function isPrime(number) that determines if the integer argument number is prime or not. The function will return a boolean True or False.

Next, write a function HowManyPrimes(P), that takes an integer P as argument and returns the number of prime numbers whose value is less than P.

And then write a function HighestPrime(K) that takes integer K as an argument and returns the highest prime that is less than or equal to K.

USE THE WHILE LOOP

Language: Python
Answer #1

Please go through the attached screenshot for indentation

Code Screenshot

def isPrime(number):
while(number<=1):
return False
while (number <=3):
return True
while (number %2 ==0 or number %3 ==0):
return False
i=5
while (i*i<=number):
while(number%i==0 or number %(i+2)==0):
return False
i =i+6
return True

def HowManyPrimes(P):
   count = 0
   i=2
   while i<P:
       if isPrime(i):
           count += 1
       i=i+1
   return count


def HighestPrime(K):
   while K>2:
       if isPrime(K):
           return K
       K=K-1

Please don't hesitate to contact me if you have any queries or require any modifications :)
Do rate the answer if it was helpful thanks.